# Defending Against Sophisticated Poisoning Attacks with RL-based Aggregation in Federated Learning

### v1

@@ Appendix D Comparison with Advanced Baselines
To demonstrate the effectiveness of AdaAggRL, we also compare it with more novel baselines, Feddefender [23] and FedVal [30], on the CIFAR-10 dataset. Using ResNet18 as the base model and a learning rate of 0.01, we conducted FL for 100 epochs. We records the accuracy, as shown in the table 3. We can see that our AdaAggRL method is still significantly better than the latest proposed baselines.
@@ Appendix E NLP Dataset Distribution Changes Caused by Poison Attacks
In order to verify whether the hypothesis that poisoning attacks can cause abnormal changes in the distribution of datasets reconstructed by gradient inversion is valid on NLP datasets, we select IMDB review dataset [19] on TextCNN [14] framework for FL semantic classification. IPM, LMP, EB and RL-attack poisoning attacks are also carried out during training. Unlike the image data set, we record and observe changes in the distribution of embedding. Figure LABEL:fig:nlp shows the statistical results of the similarity between the current client data distribution and its historical data distributions under four types of attacks that vary with the training epochs. It can be seen that on IMDB dataset, under different poisoning attacks, the similarity between the historical embedding distribution and the current embedding distribution of malicious clients is generally lower than that of benign clients. This result is consistent with the experimental results on image datasets.

### v2

@@ Appendix D Comparison with Advanced Baselines
To demonstrate the effectiveness of AdaAggRL, we also compare it with more novel baselines, Feddefender (Park et al. 2023) and FedVal (Valadi et al. 2023), on the CIFAR-10 dataset. Using ResNet18 as the base model and a learning rate of 0.01, we conducted FL for 100 epochs. We records the accuracy, as shown in the table 4. We can see that our AdaAggRL method is still significantly better than the latest proposed baselines.
@@ Appendix E NLP Dataset Distribution Changes Caused by Poison Attacks
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 13: The statistical results of the similarity between the current client data distribution and its historical data distributions under four types of attacks that vary with the training epochs.
In order to verify whether the hypothesis that poisoning attacks can cause abnormal changes in the distribution of datasets reconstructed by gradient inversion is valid on NLP datasets, we select IMDB review dataset (Maas et al. 2011) on TextCNN (Kim 2014) framework for FL semantic classification. IPM, LMP, EB and RL-attack poisoning attacks are also carried out during training. Unlike the image data set, we record and observe changes in the distribution of embedding. Figure 13 shows the statistical results of the similarity between the current client data distribution and its historical data distributions under four types of attacks that vary with the training epochs. It can be seen that on IMDB dataset, under different poisoning attacks, the similarity between the historical embedding distribution and the current embedding distribution of malicious clients is generally lower than that of benign clients. This result is consistent with the experimental results on image datasets.